OpenStudy (anonymous):

Which of the following represent the zeros of f(x) = 6x^3 – 31x^2 + 4x + 5 ?

OpenStudy (anonymous):

choices: –5, one–third, one–half 5, one–third, one–half 5, one–third, – one–half 5, –one–third, one–half

OpenStudy (anonymous):

the zeros are the x intercepts I believe

ganeshie8 (ganeshie8):

yes

OpenStudy (anonymous):

So we need to find the x intercepts lol

ganeshie8 (ganeshie8):

yes heard of 'rational root theorem' before ?

OpenStudy (anonymous):

Yeah I think so, and I got it narrowed down to two answers, the first x intercept is + 5 not -5

ganeshie8 (ganeshie8):

cool :) according to rational root theorem, possible rational zeroes will be of form :- p/q p = factors of 5 = \(\pm1, \pm 5\) q = factors of 6 \(\pm1, \pm2, \pm3, \pm6\)

OpenStudy (anonymous):

so 5, and then 1 over something and 1 over someting basically?

ganeshie8 (ganeshie8):

thats right !

OpenStudy (anonymous):

so now we need to figure out what the denominators will be

ganeshie8 (ganeshie8):

yes, there is a standard method to find the remaining zeroes, u familiar wid synthetic division also ?

ganeshie8 (ganeshie8):

since u knw that 5 is a zero, just divide (x-5), then u wil get a quadratic.

ganeshie8 (ganeshie8):

knw synthetic division ?

OpenStudy (anonymous):

Kind of know synthetic

ganeshie8 (ganeshie8):

okay, lets divide : \((6x^3 – 31x^2 + 4x + 5) \div (x-5)\)

ganeshie8 (ganeshie8):

5 | 6 -31 4 5 | ---------------------- 6

ganeshie8 (ganeshie8):

see if u can complete the rest :)

OpenStudy (anonymous):

5| 6 -31 4 5 | 30 -5 -5 -------------------- 6 -1 -1 0

ganeshie8 (ganeshie8):

Excellent ! so the depressed quadratic is ? (look at bottom row)

OpenStudy (anonymous):

6 - 1 - 1?

ganeshie8 (ganeshie8):

yes, depressed quadratic wud be :- \(6x^2 - 1x - 1\)

ganeshie8 (ganeshie8):

which u can solve using any of ur favorite method.

ganeshie8 (ganeshie8):

factor it may be...

OpenStudy (anonymous):

factoring it... -1x(5x - 1)?

OpenStudy (anonymous):

Still there?

ganeshie8 (ganeshie8):

\(6x^2 - 1x - 1\) \(6x^2 - 3x + 2x - 1\) \(3x(2x - 1) + 1(2x - 1)\) \((3x + 1)(2x - 1)\) set it to 0 x = -1/3, x = 1/2

ganeshie8 (ganeshie8):

sorry my phone rang :|

ganeshie8 (ganeshie8):

you still here ha ?

OpenStudy (anonymous):

yeah lol

OpenStudy (anonymous):

so 5, -1/3, 1/2?

ganeshie8 (ganeshie8):

yes :)